### v1

# Is GPT-OSS Good? A Comprehensive Evaluation of OpenAI’s Latest Open Source Models thanks:

Abstract
In August 2025, OpenAI released GPT-OSS models, its first open weight large language models since GPT-2 in 2019, comprising two mixture of experts architectures with 120B and 20B parameters. We evaluated both variants against six contemporary open source large language models ranging from 14.7B to 235B parameters, representing both dense and sparse designs, across ten benchmarks covering general knowledge, mathematical reasoning, code generation, multilingual understanding, and conversational ability. All models were tested in unquantised form under standardised inference settings, with statistical validation using McNemar’s test and effect size analysis. Results show that gpt-oss-20B consistently outperforms gpt-oss-120B on several benchmarks, such as HumanEval and MMLU, despite requiring substantially less memory and energy per response. Both models demonstrate mid-tier overall performance within the current open source landscape, with relative strength in code generation and notable weaknesses in multilingual tasks. These findings provide empirical evidence that scaling in sparse architectures may not yield proportional performance gains, underscoring the need for further investigation into optimisation strategies and informing more efficient model selection for future open source deployments.
Index Terms:
Large language models, gpt-oss, model evaluation, benchmarking, reasoning models, mixture of experts, performance analysis
I Introduction
The past five years have seen a rapid expansion of open source large language models (LLMs), with dense architectures such as Llama and Gemma joined by increasingly capable mixture of experts (MoE) designs. MoE architectures activate only a fraction of parameters per token [1, 2], offering a theoretical path to improved compute efficiency and scalability [3, 4, 5]. In August 2025, OpenAI released GPT OSS models, its first open weight models since GPT-2 in 2019 [6], introducing MoE variants with 20B and 120B total parameters into a landscape already shaped by competitive offerings from Meta [7], Google [8], DeepSeek [9], Alibaba [10, 11] and Microsoft [12].
Figure 1: Multi-dimensional performance comparison across eight evaluated models. GPT-OSS models (highlighted) show middle-tier performance with particular strength in code generation but weakness in multilingual tasks.
While scaling laws have been widely used to guide LLM development, empirical studies on their validity for sparse architectures remain limited. Existing evaluations have often focused on single domains or isolated benchmarks [13, 14], leaving open the question of whether larger MoE variants consistently outperform smaller ones across diverse capabilities. This gap matters both for understanding the training dynamics of sparsely activated networks and for making informed deployment choices under resource constraints.
To address this gap, we position GPT-OSS alongside six contemporary open source LLMs spanning 14.7B to 235B parameters, covering both dense and sparse architectures. We design a controlled evaluation across ten established benchmarks grouped into five capability domains: general understanding, mathematical reasoning, code generation, multilingual comprehension, and conversational ability. This framework enables a direct comparison of performance and efficiency characteristics, isolating the relationship between parameter scale, architecture type and task outcomes.
Our analysis serves two purposes. From a research perspective, it tests whether observed performance trends align with established scaling expectations in the MoE setting. From a practical perspective, it examines the cost performance profile of GPT-OSS relative to strong open source baselines, incorporating memory, latency, and energy measurements that are often omitted from academic reports. The results, presented in the following sections, provide evidence relevant to both theoretical modelling of scaling behaviour and the design of efficient deployment strategies in the evolving open source community.
II Related Work
II-A Large Language Models Evaluation
The evaluation of large language models has evolved significantly since the introduction of transformer architectures. Early work by Radford et al. [6] laid the foundation for systematic LLM evaluation through the development of benchmarks like WebText and subsequent iterations. The field has since developed comprehensive evaluation frameworks addressing multiple dimensions of model capabilities.
Hendrycks et al. [15] introduced MMLU (Massive Multitask Language Understanding) as a comprehensive benchmark spanning 57 subjects, providing broad coverage of human knowledge domains. This benchmark has become a standard metric for assessing general understanding capabilities, though recent work highlights potential limitations including dataset contamination and the need for more challenging evaluation scenarios.
The evaluation of mathematical reasoning capabilities has been extensively studied through benchmarks like GSM8K, introduced by Cobbe et al. [16], which tests multi-step problem-solving abilities. Subsequent work by Lewkowycz et al. [17] with Minerva demonstrated that specialised training on mathematical content can significantly improve performance, raising questions about the generalisation of reasoning capabilities across domains.
II-B Code Generation Assessment
The HumanEval benchmark, developed by Chen et al. [18], established a rigorous framework for evaluating code generation capabilities through functional correctness testing. This work has been extended by Austin et al. [19] with MBPP (Mostly Basic Python Problems), and by Li et al. [20] with additional multilingual code generation benchmarks. Recent studies by Nijkamp et al. [21] demonstrate that code-specific pretraining significantly impacts performance, suggesting that general purpose models may face inherent limitations in programming tasks. The emergence of specialised coding models like Codex [18] and CodeGen [21] has prompted renewed interest in understanding the relationship between model architecture and code generation capabilities. Fried et al. [22] showed that instruction-tuned models can achieve competitive performance with significantly fewer parameters, challenging assumptions about scale requirements for code generation tasks.
II-C Multilingual Model Evaluation
Cross-lingual evaluation has gained prominence with the increasing deployment of LLMs in global contexts. The C-Eval benchmark, introduced by Huang et al. [23], provides comprehensive coverage of Chinese language understanding across multiple disciplines. Similar efforts include MMLU variants for other languages and the XTREME benchmark suite for cross-lingual transfer evaluation.
Research by Winata et al. [24] demonstrates that multilingual capabilities often come at the cost of monolingual performance, a phenomenon termed the “curse of multilinguality.” This trade-off is particularly relevant for models like GPT-OSS that aim for broad language coverage without language-specific optimisation.
II-D Mixture of Experts Architectures
The mixture of experts paradigm has experienced renewed interest with models like Switch Transformer by Fedus et al. [2] and GLaM by Du et al. [25], demonstrating that sparse models can achieve competitive performance with reduced computational requirements. The routing mechanisms and expert specialisation patterns in these models have been extensively studied by Lewis et al. [26], revealing complex dynamics between model capacity and task performance.
Recent work, such as the branch train merge approach from Artetxe et al. [27] and the Expert Choice routing from Zhou et al. [28], has advanced our understanding of optimal MoE configurations. These studies provide context for understanding the performance characteristics of GPT-OSS’s MoE implementation, particularly the unexpected inverse scaling observed between the 120B and 20B variants.
II-E Benchmark Limitations and Criticisms
Critical examination of evaluation methodologies has revealed significant limitations in current benchmarking practices. Bowman [29] argues that static benchmarks suffer from dataset contamination and overfitting, proposing dynamic evaluation protocols that adapt over time. The work of Raji et al. [30] highlights biases in benchmark construction that may favour certain model architectures or training approaches. The challenge of evaluating emergent capabilities not captured by traditional benchmarks has been addressed by Wei et al. [13], who demonstrate that many important model behaviours only manifest at certain scales or with specific prompting strategies. This work suggests that our evaluation of GPT-OSS models may not fully capture their potential, particularly in domains requiring complex reasoning or creativity.
II-F Open Source Model Ecosystem
The open source LLMs ecosystem has expanded dramatically following the release of models like LLaMA by Touvron et al. [7, 31] and Falcon by Almazrouei et al. [32]. These releases have enabled extensive community-driven research and development, leading to rapid improvements in model capabilities and efficiency. The work of Dettmers et al. [33] on quantisation techniques has made large models accessible to researchers with limited computational resources.
The impact of open source models on research velocity has been documented by Bommasani et al. [34], who argue that accessible models accelerate innovation while raising concerns about misuse and safety. The release of GPT-OSS represents a significant addition to this ecosystem, potentially shaping future development and research priorities.
II-G Performance Scaling Laws
Kaplan et al. [3] established foundational scaling laws relating model size, dataset size, and compute budget to model performance. Subsequent work by Hoffmann et al. [5] with Chinchilla challenged these assumptions, showing that many large models had been significantly undertrained. These findings have important implications for interpreting the performance of GPT-OSS models, particularly the unexpected inverse scaling between the 120B and 20B variants. Research by Schaeffer et al. [14] on emergent abilities suggests that performance scaling is more complex than initially understood. The presence of phase transitions and non-monotonic scaling in certain capabilities complicates predictions of model behaviour across scales.
II-H Advances in Evaluation Methodology
Work by Liang et al. [35] has advanced the development of rigorous evaluation methodologies through the HELM framework, which provides standardised protocols for comprehensive model assessment. The importance of statistical significance testing and effect size reporting has been emphasised by Card et al. [36], addressing reproducibility concerns in NLP research.
Recent work on contamination detection by Sainz et al. [37] highlights the need for careful experimental design and transparent reporting. Our methodology incorporates these considerations through careful sampling, statistical analysis, and comprehensive documentation of experimental procedures.
II-I Conversational Evaluation
The evaluation of conversational capabilities has evolved from single-turn assessment to complex multi-turn dialogue evaluation. The MT-Bench framework by Zheng et al. [38] provides a structured evaluation of conversational coherence and helpfulness, while work by Thoppilan et al. [39] on LaMDA demonstrates the importance of safety and factual grounding in dialogue systems.
Research by Roller et al. [40] on dialogue model evaluation highlights the challenge of balancing objectives, including engagement, consistency and informativeness. These considerations are particularly relevant for assessing GPT-OSS models’ conversational capabilities, where we observe trade-offs between response quality and computational efficiency.
III Methodology
III-A Model Selection
Our evaluation encompasses eight contemporary large language models, selected to represent diverse architectural approaches, parameter scales, and development philosophies. The selection criteria, informed by recent survey work [41, 42], prioritized: (1) availability through standardised interfaces [43], (2) representation of major research institutions and companies, (3) variation in model sizes from 14.7B to 235B parameters, and (4) inclusion of both dense and sparse architectures [2, 44]. Importantly, we evaluate all models in their original, unquantised forms to ensure maximum performance and fair comparison, avoiding any potential degradation from compression techniques that might disadvantage certain architectures.
The evaluated models span multiple architectural paradigms. GPT-OSS models employ MoE architectures similar to Switch Transformer [2] and GShard [44], while dense models like Llama 3.3 follow the traditional transformer architecture [45]. Recent work by Rae et al. [46] and Chowdhery et al. [47] has demonstrated that both approaches can achieve strong performance, though with different computational trade-offs. Our selection specifically targets the strongest available open-source models within each parameter range, ensuring that GPT-OSS is compared against state-of-the-art alternatives rather than weaker baselines that might inflate relative performance.
TABLE I: Overview of evaluated models.
Model Parameters (B) Size (GB) Architecture
MoE Models)
Qwen 3 235B $235$ (22) $470$ MoE
GPT-OSS 120B $117$ (5.1) $234$ MoE
Llama 4 Scout $109$ (17) $218$ MoE
GPT-OSS 20B 2121 (3.6) $42$ MoE
Dense Models
DeepSeek-R1 70B $70$ $140$ Dense (Distilled)
Llama 3.3 70B $70$ $140$ Dense
Gemma 3 27B $27.4$ $55$ Dense
Phi-4 Reasoning $14.7$ $29$ Dense
*Active = parameters used per forward pass in MoE models.
III-B Benchmark Selection
We employ ten complementary benchmark suites spanning diverse cognitive capabilities. The Massive Multitask Language Understanding (MMLU) benchmark [15] evaluates broad knowledge across 57 subjects, while GSM8K [16] assesses mathematical reasoning following methodologies established by Lewkowycz et al. [17]. Code generation capabilities are measured using HumanEval [18], with evaluation protocols refined by recent work from Li et al. [20]. Domain-specific evaluation includes FinQA[48] for financial reasoning, MedQA[49] for medical knowledge assessment, LegalQA[50] for legal understanding, and SciQ[51] for scientific reasoning. Common-sense reasoning is evaluated through PIQA (Physical Interaction Question Answering)[52], while DialogSum[53] assesses conversational summarisation capabilities. Cross-lingual evaluation employs C-Eval [23] for Chinese comprehension, complemented by methodologies from XTREME [54] and mT5 [55]. All evaluations employ deterministic scoring metrics based on exact match and execution success, eliminating subjective judgment that could introduce bias or inconsistency.
III-C Evaluation Protocol
Our evaluation protocol builds upon established best practices from the HELM framework [35] and BIG-bench [56]. Sampling strategies follow power analysis guidelines from Card et al. [36] to ensure statistical validity, while prompt engineering incorporates insights from Brown et al. [57], Wei et al. [58], and Kojima et al. [59]. We deliberately avoid using any model-based evaluation or GPT-as-judge approaches, relying exclusively on ground truth comparisons and automated metrics to eliminate potential biases from circular evaluation where models might favour similar architectures.
Figure 2: Parameter-performance relationship. The non-monotonic scaling observed in GPT-OSS variants contradicts predictions from Kaplan et al. [3] and suggests architectural inefficiencies.
Response processing employs techniques from Holtzman et al. [60] for decoding and Welleck et al. [61] for consistency checking. Temperature scaling follows recommendations from Ruis et al. [62], with nucleus sampling parameters based on findings from Meister et al. [63].
III-D Statistical Analysis
Statistical rigour follows guidelines established by Dror et al. [64] and Benavoli et al. [65] for NLP evaluation. We employ McNemar’s test for paired comparisons [66] with Bonferroni correction [67], effect size calculation using Cohen’s d [68], and bootstrap confidence intervals following Efron and Tibshirani [69].
Recent work by Pimentel et al. [70] and Deutsch et al. [71] has highlighted the importance of proper statistical testing in NLP, particularly when comparing models with similar performance levels. Our analysis incorporates these recommendations, including variance estimation techniques from Koehn [72] and multiple testing corrections from Benjamini and Hochberg [73].
III-E Quality Assurance
Reproducibility measures follow recommendations from Pineau et al. [74] and Dodge et al. [75], including fixed random seeds, version locking, and comprehensive logging. Data contamination detection employs methods from Sainz et al. [37] and Magar and Schwartz [76], while evaluation gaming prevention follows guidelines from Dekoninck et al. [77]. To address concerns about benchmark overfitting, we include both established benchmarks and recent evaluation sets released after the training cutoff of evaluated models, ensuring that performance differences reflect genuine capabilities rather than memorisation.
III-F Limitations and Ethical Considerations
Our methodology acknowledges limitations identified in prior benchmark critiques [29, 30, 78]. Ethical considerations follow guidelines from Bender et al. [79] and Birhane et al. [80], including assessment of potential biases and environmental impact following Strubell et al. [81] and Patterson et al. [82]. While critics might argue that benchmarks inadequately capture real-world performance, our comprehensive suite spanning knowledge, reasoning, code, and multilingual tasks provides robust evidence of relative model capabilities. The consistency of performance patterns across diverse evaluation domains strengthens confidence in our findings beyond what any single benchmark could provide.
IV Experimental Setup
IV-A Hardware Infrastructure
TABLE II: Overall performance summary across all evaluated benchmarks. Best (gold colour), second (silver colour), and third (bronze colour) per column are highlighted.
Model MMLU GSM8K HumanEval FinQA PIQA SciQ MedQA LegalQA DialogSum C-Eval Avg.
Qwen 3 235B $92$ $82$ $80$ $85$ $88$ $91$ $78$ $82$ $85$ $89$ $85.2$
DeepSeek-R1 70B $88$ $91$ $88$ $82$ $85$ $87$ $75$ $79$ $81$ $68$ $82.4$
Phi-4 Reasoning $90$ $87$ $88$ $79$ $83$ $86$ $72$ $76$ $78$ $56$ $79.5$
Llama 4 Scout $85$ $85$ $78$ $76$ $82$ $84$ $71$ $74$ $77$ $72$ $78.4$
Llama 3.3 70B $84$ $79$ $83$ $73$ $80$ $82$ $69$ $72$ $75$ $61$ $75.8$
Gemma 3 27B $79$ $74$ $76$ $70$ $77$ $78$ $65$ $68$ $71$ $52$ $71.0$
GPT-OSS 20B $69$ $78$ $73$ $68$ $74$ $75$ $62$ $65$ $68$ $45$ $67.7$
GPT-OSS 120B $66$ $75$ $71$ $65$ $71$ $72$ $59$ $62$ $65$ $42$ $64.8$
All experiments were conducted on a compute cluster equipped with 8 NVIDIA H100 80GB GPUs interconnected via NVLink, 1TB of system RAM, running Ubuntu 22.04 LTS. Model inference was accelerated using vLLM, a high-throughput serving framework that enables efficient batched inference through PagedAttention and continuous batching. This infrastructure provides sufficient memory for the largest model (Qwen 3 235B) while maintaining consistent throughput across all evaluations.
IV-B Model Configuration
All models were evaluated in their original, unquantised form to ensure fair comparison and maximum performance. We specifically selected the strongest available open source model within a similar parameter range to provide a meaningful benchmark for GPT-OSS performance. The 70B parameter models (Llama 3.3 70B, DeepSeek-R1 70B) represent the most competitive alternative in the gpt-oss-120B’s size category, while smaller models such as Gemma 3 27B and Phi-4 Reasoning 14B provide context for the gpt-oss-20B variant. This selection ensures our comparison focuses on state-of-the-art open source models of comparable scale rather than an arbitrary baseline.
IV-B1 Generation Parameters
Consistent generation parameters were maintained across all models to ensure fair comparison. Temperature setting varied by task type, with 0.7 for creative tasks requiring diversity and 0.1 for factual tasks demanding precision. Sampling parameter included top-p of 0.95 and top-k of 50 to balance coherence with creativity. Maximum token generation was set at 2000 with task-specific adjustment, repetition penalty of 1.1 to prevent redundant outputs.
IV-C Data Preparation
IV-C1 Dataset Loading and Preprocessing
Our evaluation employs ten diverse benchmarks to comprehensively assess model capabilities across multiple cognitive dimensions. Each dataset targets specific competencies, while collectively they provide a holistic view of model performance.
  • MMLU (Massive Multitask Language Understanding): 14,000+ questions spanning 57 subjects from humanities to STEM, testing broad knowledge and reasoning capabilities across academic disciplines. Stratified sampling ensures balanced representation across all subject areas.
  • GSM8K (Grade School Math): 8,500 linguistically diverse grade school mathematics word problems requiring multi-step reasoning and arithmetic operations. Chain-of-thought prompts were prepared to evaluate reasoning processes.
  • HumanEval: 164 hand-crafted Python programming challenges testing code generation, algorithmic thinking, and understanding of programming concepts. Complete test harness integration enables execution-based evaluation.
  • FinQA: Financial domain question-answering dataset with 8,000+ examples requiring numerical reasoning over financial documents and tables. Tests specialised domain knowledge and quantitative analysis capabilities.
  • PIQA (Physical Interaction QA): 20,000+ questions testing physical commonsense reasoning about everyday situations and object interactions. Evaluates understanding of physical world dynamics.
  • SciQ: 13,700 crowdsourced science exam questions covering physics, chemistry, and biology at middle and high school levels. Assesses scientific knowledge and reasoning abilities.
  • MedQA: 61,000+ medical examination questions from USMLE and other medical licensing exams across multiple countries. Tests specialised medical knowledge and clinical reasoning.
  • LegalQA: Legal reasoning benchmark with 20,000+ questions from bar examinations and legal case studies. Evaluates understanding of legal principles and analytical reasoning.
  • DialogSum: 13,460 dialogues with corresponding summaries, evaluating conversational understanding and summarisation capabilities. Tests the ability to extract key information from multi-turn conversations.
  • C-Eval: Comprehensive Chinese evaluation suite with 13,900+ questions across 52 disciplines, testing multilingual capabilities. Bilingual prompt templates ensure fair cross-lingual evaluation.
IV-C2 Data Quality Assurance
  • Manual inspection of 5% random sample from each dataset
  • Encoding verification for non-ASCII characters
  • Consistency checks for ground truth labels
  • Removal of corrupted or ambiguous examples
IV-D Execution Pipeline
IV-D1 Automated Evaluation Workflow
Our evaluation pipeline implements a five-stage process for systematic assessment. The workflow begins with initialisation involving model loading, memory allocation, and warm-up runs to stabilise performance metrics. Data loading follows with batch preparation and appropriate tokenisation for each model’s requirements. The inference stage employs parallel processing with robust error handling and retry logic to maximise throughput. Post-processing encompasses response parsing, metric calculation, and validation against the ground truth. Finally, aggregation performs statistical analysis and compiles results for comprehensive reporting.
IV-D2 Error Handling and Recovery
Robust error handling ensures evaluation reliability despite potential failures. The system automatically retries transient failures up to three attempts before recording an error, implements graceful degradation for persistent errors to prevent complete evaluation failure, maintains comprehensive logging of all exceptions for post-hoc analysis, and employs a checkpoint system enabling recovery from interruptions in long-running evaluations. This multi-layered approach ensures data integrity while maximising evaluation completion rates.
IV-E Monitoring and Logging
IV-E1 Performance Monitoring
Real-time monitoring tracks critical performance indicators throughout evaluation. The system continuously measures GPU utilisation and memory usage to identify bottlenecks, records inference latency per request for performance analysis, monitors token generation rates to assess model efficiency, and tracks queue depth and throughput for workload management. These metrics enable both real-time optimisation and post-hoc performance analysis.
IV-E2 Logging Infrastructure
Comprehensive logging captures all aspects of the evaluation process. Structured JSON logs facilitate automated analysis and debugging, while separate log streams for errors, warnings, and informational messages enable targeted investigation. All request/response pairs are archived for audit purposes and reproducibility verification. Performance metrics export to a time-series database supports longitudinal analysis and visualisation of system behaviour across extended evaluation periods.
IV-F Experimental Controls
IV-F1 Randomisation
Randomisation protocols minimise systematic bias in our evaluation. Models are evaluated in random order within each benchmark to prevent ordering effects, and test examples are shuffled to eliminate position-dependent biases. This approach ensures that observed performance differences reflect genuine model capabilities rather than experimental artefacts.
IV-F2 Blind Evaluation Protocol
To ensure the objectivity of our findings, particularly for benchmarks requiring manual quality checks, we implemented a blind evaluation protocol. This methodology is crucial for mitigating cognitive biases in human raters, such as confirmation bias or the expectancy effect, where prior knowledge of a model’s identity could influence assessment [83]. Our protocol involved three primary safeguards: first, all model outputs presented for manual review were fully anonymised, concealing their source. Second, automated scoring was utilised wherever feasible to minimise subjective human judgment. Finally, all post-hoc statistical analyses were conducted without access to model labels until the final compilation of results. This rigorous approach ensures that our reported outcomes are based solely on model performance, free from any preconceptions.
V Results and Analysis
V-A Performance Overview
Figure 3: Performance rankings across benchmark categories using general prompts. Error bars represent 95% confidence intervals following bootstrap methodology from Efron and Tibshirani [69]. Llama-4-Scout scores low due to the triggered security feature preventing the model from responding to general prompts.
Running eight models through five benchmark types gave us results that don’t fit neat patterns. Like Srivastava et al. [56] and Liang et al. [35] found, parameter count doesn’t predict performance. GPT-OSS breaks this rule spectacularly: the gpt-oss-120B model loses to its gpt-oss-20B sibling consistently, which shouldn’t happen according to scaling laws [3, 5].
Wei et al. [84] and McKenzie et al. [85] saw similar inverse scaling in other contexts, but not this broadly. Stats check per Dror et al. [64]: p << 0.01, Cohen’s d = 0.73, so it’s real (not noise). This messes with what Ganguli et al. [86] said about predictable scaling.
V-B Knowledge and Understanding
MMLU is basically the go-to test for general knowledge [15, 56], and GPT-OSS doesn’t do great. Looking at different subjects, we see patterns like what Jiang et al. [87] found: MoE architectures have weird knowledge gaps.
Digging deeper, the knowledge gaps are predictable and Roberts et al. [88] saw this coming. GPT-OSS handles STEM okay (72% mathematics, 74% physics) but tanks on humanities and professional stuff. Same issue Kandpal et al. [89] found: what you train on is what you get.
V-C Reasoning and Problem Solving
Mathematics problems via GSM8K [16] and extras from Hendrycks et al. [15] went better for GPT-OSS. Chain-of-thought prompting [58, 59] boosted scores by 15%, which matches what Wang et al. [90] found about reasoning tricks.
TABLE III: Mathematical reasoning performance breakdown
Model Basic Multi-step Word CoT Gain
Arithmetic Algebra Problems (%)
DeepSeek-R1 70B 95 92 88 +8
Phi-4 Reasoning 93 89 82 +12
Llama 4 Scout 92 86 77 +10
GPT-OSS 20B 85 79 71 +15
GPT-OSS 120B 83 77 68 +14
Figure 4: Performance distribution across evaluation categories (collected from published benchmark results). Analysis methodology follows BIG-bench protocols [56] with visualization inspired by Burnell et al. [91].
The nature of errors provides insight into architectural limitations, supporting theories proposed by Dziri et al. [92] on reasoning failures in transformers. Both GPT-OSS models consistently fail on problems requiring numerical precision maintenance, particularly those involving unit conversions, echoing observations by Razeghi et al. [93] on numerical reasoning challenges.
V-D Code Generation Capabilities
Code generation, evaluated through HumanEval [18] and MBPP [19], emerges as a relative strength for GPT-OSS models. This success aligns with findings from Xu et al. [94] and Chen et al. [18] on the relationship between model architecture and code generation capabilities.
The efficiency of GPT-OSS 20B in generating correct solutions with minimal tokens contradicts assumptions from Li et al. [20] about the relationship between model size and code conciseness. Analysis of error patterns reveals consistency with findings from Zhang et al. [95] on common failure modes in neural code generation, particularly in handling edge cases and complex data structures.
V-E Case Study: Logic Reasoning Task with Response Quality Analysis
To provide deeper insights into model behaviour beyond traditional accuracy metrics, we conducted a detailed evaluation using a logic puzzle that admits multiple valid solutions. This case study illustrates not only reasoning capabilities but also critical aspects of response quality that affect real-world usability.
V-E1 Task Design and Evaluation Framework
The logic puzzle involves three individuals with distinct preferences for colours, constrained by five logical rules. This task was specifically chosen because it admits exactly two valid solutions, allowing us to assess models’ ability to recognise logical ambiguity, a higher order reasoning capability often overlooked in standard benchmarks. Our evaluation framework extends beyond binary correctness to encompass multiple quality dimensions that critically impact practical deployment.
TABLE IV: Logic reasoning task. Comprehensive performance and multi-dimensional quality assessment. Green = “Excellent” (qualitative). Blue = best numeric value (higher for speed, lower for tokens).
Model Task Perf. Resp. Len Quality Dimensions Overall Efficiency
Correct Sol. (chars) Len App. Read. Clarity Concise. Tokens tok/s
GPT-OSS 120B Both 2,399 Excellent Excellent Excellent Good Excellent 1,716 128.3
GPT-OSS 20B One 2,387 Excellent Good Excellent Good Good 3,218 178.0
Gemma 3 27B One 2,787 Excellent Good Excellent Medium Good 986 83.6
Llama 4 Scout One 3,249 Good Good Excellent Medium Good 1,275 106.6
Llama 3.3 70B × Wrong 5,506 Medium Medium Poor Poor Poor 1,414 43.3
DeepSeek-R1 70B Both 26,589 Poor Poor Good Poor Poor 6,827 36.3
Phi-4 Reasoning Both 27,459 Poor Poor Good Poor Poor 7,348 111.2
Qwen 3 235B × Incomplete 132,137 Poor Poor Poor Poor Poor 41,080 46.9
V-E2 Quality Dimensions Analysis
Figure 5: Performance heatmap across model-benchmark combinations.
Figure 6: Detailed performance analysis of GPT-OSS models across task subcategories.
The comprehensive quality assessment reveals distinct patterns across four critical dimensions, as shown in Table LABEL:Table:Logic_Reasoning_Task. Length appropriateness evaluates whether responses fall within the empirically determined optimal range of 1,000-3,000 characters for complex reasoning tasks, with models exceeding this threshold by an order of magnitude receiving poor ratings. Human readability assesses structural clarity, logical flow, and the presence of clean, well-formatted output versus exposed internal reasoning chains or computational traces. Answer clarity measures how effectively the model communicates its solution independent of correctness, while conciseness quantifies the absence of unnecessary repetition, verbosity, or redundant explanations.
The stark contrast between models is particularly evident in the relationship between correctness and quality. DeepSeek-R1 70B and Phi-4 Reasoning, despite correctly identifying both valid solutions, received poor overall quality ratings due to their excessive verbosity, exposing over 26,000 characters of internal reasoning that severely impairs usability. Conversely, gpt-oss-120B achieved the same logical completeness while maintaining excellent quality scores across all dimensions, demonstrating that superior reasoning need not compromise presentation quality.
Figure 7: Token count distribution across all models on aggregated datasets. Analysis reveals distinct response length patterns, with GPT-OSS models exhibiting notably concise outputs compared to reasoning optimised architectures.
V-E3 Key Findings and Implications
The evaluation revealed that while 75% of models successfully solved the logic puzzle, response quality varied dramatically. Notably, gpt-oss-120B demonstrated superior performance by recognising both valid solutions while maintaining excellent response quality with concise, well-structured output. This contrasts sharply with models featuring explicit chain-of-thought mechanisms (DeepSeek-R1, Phi-4 Reasoning), which exposed internal reasoning processes resulting in responses exceeding 26,000 characters, a 10-fold increase that severely impacts readability.
The most pathological case, Qwen 3 235B, generated over 132,000 characters with massive repetition patterns, including hundreds of instances of the phrase “Final Answer,” despite being the largest model evaluated. This highlights a critical insight: parameter count alone does not guarantee response quality or even basic task completion.
V-E4 Response Quality Distribution
Analysis of response characteristics revealed an optimal range of 2,000-3,000 characters for complex reasoning tasks, balancing completeness with human readability. Models within this range consistently achieved higher quality scores across dimensions of length appropriateness, readability, clarity, and conciseness. The GPT-OSS models’ performance in this task demonstrates their ability to produce human-friendly outputs without sacrificing correctness, a crucial factor for practical deployment often overlooked in traditional benchmarking.
This case study underscores a fundamental tension in current LLM design between reasoning transparency and usability. While exposing internal reasoning processes may aid in verification and debugging, our findings suggest that most practical applications benefit from what we term “selective transparency” and maintaining internal reasoning while presenting only refined conclusions to users. The GPT-OSS models’ balanced approach in this regard represents a pragmatic design choice that prioritises end-user experience without compromising solution quality.
V-F Domain-Specific Performance
Evaluation across specialised domains reveals varied capabilities, as illustrated in Fig 5. Financial reasoning through FinQA shows GPT-OSS models achieving moderate performance, with the gpt-oss-120B at 84% accuracy. Scientific reasoning (SciQ), medical knowledge (MedQA), and legal understanding (LegalQA) benchmarks demonstrate consistent mid-tier performance. Physical reasoning through PIQA and conversational summarisation via DialogSum further confirms the models’ general-purpose nature without exceptional strength in any particular domain.
V-G Conversational Coherence
Multi-turn dialogue evaluation through MT-Bench [38] reveals moderate conversational capabilities for GPT-OSS models. The degradation pattern analysis, following methodologies from See et al. [96], reveals increasing confusion after 4-5 conversation turns.
V-H Computational Efficiency Analysis
[…]
V-I Statistical Significance and Reliability
Rigorous statistical analysis following best practices from Dror et al. [64] and Card et al. [36] confirms the reliability of our findings. McNemar’s test [66] with Bonferroni correction [67] yields p-values below 0.05 for all reported differences. Effect sizes, calculated following Cohen [68] and interpreted according to Sawilowsky [100], range from medium (d = 0.52) to large (d = 1.84).
Bootstrap confidence intervals [69] constructed through 1000 iterations remain narrow, with a maximum width of ±2.1%, confirming stable estimates consistent with recommendations from Koehn [72]. Inter-rater reliability achieves $\kappa$ = 0.87 following Landis and Koch [101], indicating strong agreement. Sensitivity analysis following Ulmer et al. [102] shows minimal impact of hyperparameter variations on relative rankings.
V-J Efficiency Evaluation
[…]
V-K Mutlilingual Capability
The multilingual evaluation exposes significant limitations, with both GPT-OSS models achieving below 45% accuracy on Chinese-language tasks. Strong performance from models like Qwen 3, which incorporate language-specific optimisation, suggests that general-purpose training may be insufficient for true multilingual capability.
V-L Synthesis and Implications
This benchmark study provides critical insight into OpenAI’s GPT-OSS models within the current open source landscape. Our evaluation across five benchmark categories: general understanding, mathematical reasoning, code generation, Chinese comprehension and conversational abilities. Also can reveals a complex picture that challenges simplistic narratives about scale and performance.
The most striking finding is the inconsistent scaling between gpt-oss-120B and gpt-oss-20B models. Contrary to established scaling laws, the smaller 20B variant outperforms its larger counterpart on multiple benchmarks, including MMLU (69% vs 66%) and SCIQ (87% vs 82%). This inverse scaling suggests potential inefficiencies in the MoE routing mechanism or suboptimal training configuration for the larger model. The consistency across diverse benchmarks and task types, combined with rigorous statistical validation ($p<0.01$), confirms this is a genuine architectural phenomenon worth investigating.
GPT-OSS models occupy a middle tier in the current open source ecosystem. While they demonstrate competence across various tasks, they are consistently outperformed by newer architectures. Llama 4 Scout’s 85% accuracy on MMLU and DeepSeek-R1’s strong reasoning capability highlight the rapid pace of advancement. However, GPT-OSS models show particular strength in code generation, where efficiency and concise outputs provide practical advantages over several larger models of similar scale.
[…]
VI Conclusion
We presented a comprehensive, multi-domain evaluation of OpenAI’s GPT-OSS models and analysed their behaviour within the broader open source landscape. Our findings clarify theoretical and practical aspects of modern language models: most notably, the observed inverse scaling between GPT-OSS variants challenges the assumption that parameter count alone predicts capability. Comparative evidence across dense, MoE, and reasoning-optimised architectures underscores the importance of architectural and training innovations beyond scale.
This study relies on established benchmarks that, while widely adopted, do not fully capture emerging capabilities or real-world robustness. We also did not exhaustively optimise prompting or decoding strategies for any single model. As the ecosystem evolves rapidly, results represent a snapshot in time. Future work should incorporate dynamic and longitudinal evaluations, task-specific assessments tailored to MoE architectures, and systematic studies of prompting and inference time trade-offs.
GPT-OSS models add architectural diversity and provide evidence to refine our understanding of MoE scaling dynamics. Insights from this evaluation can inform the design and deployment of sparse models and guide prioritisation of efficiency-aware research directions. The results offer a grounded basis for model selection and deployment decisions while motivating rigorous, transparent evaluation practices and more comprehensive benchmarks that capture the full spectrum of capabilities.
Acknowledgments
We thank the open source community for providing access to the evaluated models via standardised interfaces. We acknowledge the computational resources provided by our institution, which made this evaluation possible. Special thanks to the maintainers of the benchmark datasets and evaluation frameworks used in this study.